# A Bayesian workflow for securitizing casualty insurance risk

### v2

@@ 2.3 Model comparison, selection and averaging
3 Ledger’s Bayesian modeling workflow
@@
Figure 1: A Bayesian workflow for modeling casualty insurance loss development and forecasting future losses – the first step for securitization. In-depth explanations for each step are described throughout the section 3. Here, $\theta$ represents all parameters in the model, and $p(y,\theta)$ then represents the joint distribution of data and model parameters. Graphs in the workflow steps here are only included for illustration; we dive into each workflow step in more detail in the following sections.
@@ 3.1.1 Body development
In most time-series like models, we expect that uncertainty would increase across time – for loss development models, uncertainty propegation is a bit more complex. Intuitively, the further into the future we are when looking back at losses for a given accident year, the more claims are closed, and as a result we are more certain in the ultimate loss. To capture this behavior, we assume that variance in the ultimate loss follows a linear model with an intercept and slope captured by $\bm{\gamma}=(\gamma_{1},\gamma_{2})^{\prime}$. We expect $\gamma_{2}<0$, indicating that losses vary around $\mu_{ij}$ less as development progresses. This decrease in variability is critical to capture the asymptotic behavior of losses across development – without this heterogeneity mechanism, the model predictions will be too certain for early development lags and too uncertain for later development lags.
@@
Due to the two-stage nature of our loss development workflow, the chain-ladder model is only fitted to losses where $j\leq\tau$, where $\tau\in{2,...,M}$ is an integer chosen by an analyst based on their knowledge of how long the body process should be active. Generally, $\tau$ should be lower if the program is quick to develop (reaching the tail process for lower develoment lags), and higher if the program is slow to develop.
@@ 3.2 Future ultimate loss forecasting
Before defining the models, we will introduce new notation. Although we find it useful for loss development to be directly on the losses, we find it easier to work with loss ratios for the forecasting step. Below, we use $r$ to indicate the loss ratio, such that $r_{ij}=\frac{y_{ij}}{p_{i}}\forall i=1,...,N,\forall j=1,...,M$, where $p_{i}$ is the premium volume for accident year $i$.
@@ 3.2.1 A basic random walk model
where $\bm{\eta}=(\eta_{1},\dots,\eta_{N})^{\prime}$ is a latent (in log-space) ultimate loss ratio that drifts across accident years with deviations between years proportional to $\epsilon$. The “observed” ultimate loss ratio ($r_{i\infty}$) then follows a lognormal distribution with observation noise $\sigma_{i}$ that is a function of earned premium for the given accident year ($p_{i}$) such that as earned premium increases, observation noise decreases.
@@ 3.2.3 Adding measurement error
The forecasting models above are defined with $r_{i\infty}$ as the target. However, in real-world scenarios, we do not know $r_{i\infty}$ – instead, we have the posterior predictions from the loss development models, or $\tilde{y}_{i\infty}$ (transformed to the loss ratio scale per $\tilde{r}_{i\infty}=\frac{\tilde{y}_{i\infty}}{p_{i}}\forall i=1,...,N$). To account for uncertainty in these posterior predictions in the context of the forecasting model, we can add an error-in-variables measurement error model (Stefanski, 2000) to the forecasting models that takes the posterior means $\mathbb{E}\left[\tilde{r}_{i\infty}\right]\forall i=1,...,N$ and posterior standard deviations $\mathbb{SD}\left[\tilde{r}_{i\infty}\right]\forall i=1,...,N$ as input and models the “true” underlying ultimate loss ratio ${r^{\prime}}_{i\infty}$. For both the random walk and mean reversion models, this involves the following modification to the likelihood expression:
@@
In practice, there are many potential measurement error assumptions one can make, and the above is just one example. In the current context, lognormal measurement error is reasonable given that the development model predictions are generated from a lognormal distribution. An alternative approach would be to jointly model the loss development and forecasting stages such that there is no need for an intermediate summarization step and subsequent measurement error modeling (akin to Goold (2024), who jointly models the body and tail stage of loss development). However, in practice we find that it is useful to separate the loss development from forecasting stages, which allows for more intuitive setting of priors on the individual models and easier investigation of model behavior. Further, having separate stages keeps the flow of information forward such that the forecasting model does not influence development model parameter estimation.
@@ 3.3 Deriving data-driven priors
3.4 Model validation
@@ 3.4.2 Prior and posterior predictive checks
Figure 3: Prior predictions for three random commercial auto programs from the Meyers (2015) dataset. Lines are 30 different realizations from the prior distribution, and points are true losses at $y_{i(10)}$.
@@
Figure 4: Posterior predictions for three random commercial auto programs from the Meyers (2015) dataset. The programs here are the same as in Figure 3. Lines are 30 different realizations from the posterior distribution, and points are true losses at $y_{i(10)}$.
3.5 Backtesting workflow
A model can pass all model validation checks but then fail to perform well on real-world prediction tasks given the wide range of different loss triangles that we encounter from day-to-day. Therefore, large-scale out-of-sample tests are crucial to determine if a given model (or model pipeline) is suited for production use. Further, we often have multiple competing models, where each model makes different assumptions regarding how losses change over time. For example, equation 7 assumes mean reversion toward some latent ultimate loss, whereas 6 assumes no such latent ultimate loss or reversion process. An added benefit of conducting out-of-sample tests is that we can take advantage of modern Bayesian stacking methods that combine predictions from competing models, allowing for a combined performance better than the best fitting individual model. In the following sections, we overview our out-of-sample backtesting workflow and our use of stacking to improve model performance.
@@ 3.5 Backtesting workflow
For demonstration purposes, here we will rely on an open dataset described by Meyers (2015), which contains about 50 loss triangles for each of four different lines of business, including private passenger auto, worker’s compensation, commercial auto, and general liability lines. Each triangle in the Meyers dataset is 10 accident years by 10 development lags, which allows for us to fit our loss development to loss forecasting pipeline on each triangle exactly once, using the loss from the most recent accident year, development lag pair as the future, left-out datapoint for testing purposes (see Figure 5). Note that this scheme assumes $y_{i\infty}=y_{i(10)}\forall i=1,...,N$, which may not be true in practice for longer-tailed lines of business. However, given data constraints for this demonstration, we believe this is a reasonable assumption. We use $y_{i\infty}$ and $y_{i(10)}$ interchangeably in sections that follow.
@@ 3.5.1 Stacking of predictive distributions
For demonstration purposes, we will use the stacking model described by Yao et al. (2018), which estimates a single weight for each model. The model can be easily fit using bayesblend – an opensource Python library that we have developed that implements various different stacking models (Haines and Goold, 2024). To fit the stacking model, we first computed the posterior log likelihood values and posterior predictions for each datapoint in the test set as illustrated in Figure 5. We did so for both the random walk and mean reversion forecasting models. Note that each posterior log likelihood in the test set is derived based on the true loss, which the models are not trained on. We then used the posterior log likelihoods and associated posterior predictions for each model as input to the MleStacking class in bayesblend, which outputs a blend of the posterior log likelihoods and predictions given the estimated model weights.
@@ 3.5.3 Model performance results
Figure 6: ELPD (left) and RMSE (right) differences (+/- 2 standard errors on the difference), ordered by Validation dataset performance for each model and line of business. The best-performing model is the reference model, shown at the top of each panel. Absolute ELPD and RMSE values are displayed above. Positive ELPD differences with an uncertainty interval that does not cross zero indicates a credible difference at the 95% level in favour of the top model. Negative RMSE differences with an uncertainty interval that does not cross zero indicates a credible difference in favor of the top model.
@@ 3.6 From Forecasts to Cashflows
Once we have selected a model configuration that can accurately forecast ultimate losses for future accident years, the next step is to estimate cashflows that can be used to understand the profitability of the insurance program. There are two types of cashflows that are relevant for this step: (1) the “underwriting” cashflow, which captures primarily how losses are paid out across time for the future accident year being securitized; and (2) the “investment” cashflow, which captures interest on invested premium dollars (i.e. the “float”) as well as bespoke deal structure terms (collateralization schemes, profit commission to cedants, etc.). For the current demonstration, we will focus primarily on the underwriting cashflow, as the investment cashflow is a deterministic function of underwriting cashflow (i.e. it mostly an exercise in accounting) so is outside the scope the current work.
@@ 3.6.1 Estimating future underwriting cashflows
$\begin{split}\tilde{y}_{i(9)}^{(s)}&=\tilde{y}^{(s)}_{i(10)}/\alpha^{(s)}_{9}\\ \tilde{y}_{i(8)}^{(s)}&=\tilde{y}^{(s)}_{i(9)}/\alpha^{(s)}_{8}\quad\forall s=1,...,S\\ &\vdots\\ \tilde{y}_{i(1)}^{(s)}&=\tilde{y}^{(s)}_{i(2)}/\alpha^{(s)}_{1}\end{split}$ (16)
Figure 8: Underwriting cashflows for future accident years based on historic development patters and forecasted ultimate losses for a random selection of commercial auto programs. Underwriting cashflows were estimated per the walk-back method described in equation 16 using random programs selected from the Backtest workflow (section 3.5) illustration. Forecasted ultimate losses were based on the stacked predictions from the random walk 6 and mean reversion 7 models, and age-to-age factors were obtained from the chain-ladder 4 and generalized Bondy 5 models fitted to each program.
where $\alpha^{(s)}_{j}$ is the $j$th age-to-age factor and $s$ indicates the posterior sample. The result is a full posterior distribution of paid loss estimates for each future accident year $i$ (see Figure 8). Across development, these paid loss estimates, relative to premiums, correspond to the underwriting cashflow associated with the future accident year.

### v3

@@ 2.3 Model comparison, selection and averaging
3 A Bayesian workflow for casualty ILS
@@
Figure 1: A Bayesian workflow for modeling casualty insurance loss development and forecasting future losses – the first step for securitization. In-depth explanations for each step are described throughout Section 3. Here, $\theta$ represents all parameters in the model, and $p(y,\theta)$ then represents the joint distribution of data and model parameters. Graphs in the workflow steps here are only included for illustration; we dive into each workflow step in more detail in the following sections.
@@ 3.1.1 Body development
In most ARIMA or state-space style time-series models, we expect that uncertainty would increase across time – for loss development models, uncertainty propagation is a bit more complex. Intuitively, the further into the future we are when looking back at losses for a given accident year, the more claims are closed, and as a result we are more certain in the ultimate loss. To capture this behavior, we assume that variance in the ultimate loss follows a linear model with an intercept and slope captured by $\bm{\gamma}=(\gamma_{1},\gamma_{2})^{\prime}$. We expect $\gamma_{2}<0$, indicating that losses vary around $\mu_{ij}$ less as development progresses. This decrease in variability is critical to capture the asymptotic behavior of losses across development – without this heterogeneity mechanism, the model predictions will be too certain for early development lags and too uncertain for later development lags.
@@
Due to the two-stage nature of our loss development workflow, the chain-ladder model is only fitted to losses where $j\leq\tau$, where $\tau\in{2,...,M}$ is an integer chosen by an analyst based on their knowledge of how long the body development process should be active. Generally, $\tau$ should be lower if the program is quick to develop (reaching the tail process for lower development lags), and higher if the program is slow to develop.
@@ 3.2 Future ultimate loss forecasting
Before defining the models, we will introduce new notation. Although we find it useful for loss development to be directly on the losses, we find it easier to work with loss ratios for the forecasting step. Below, we use $r$ to indicate the loss ratio, such that $r_{ij}=\frac{y_{ij}}{p_{i}}~{}\forall i=1,...,N,\forall j=1,...,M$, where $p_{i}$ is the premium volume for accident year $i$.
@@ 3.2.1 A basic random walk model
where $\bm{\eta}=(\eta_{1},\dots,\eta_{N})^{\prime}$ is a latent (in log-scale) ultimate loss ratio that drifts across accident years with deviations between years proportional to $\epsilon$. The “observed” ultimate loss ratio ($r_{i\infty}$) then follows a lognormal distribution with observation noise $\sigma_{i}$ that is a function of earned premium for the given accident year ($p_{i}$) such that as earned premium increases, observation noise decreases.
@@ 3.2.3 Adding measurement error
The forecasting models above are defined with $r_{i\infty}$ as the target. However, in real-world scenarios, we do not know $r_{i\infty}$ – instead, we have the posterior predictions from the loss development models, or $\tilde{y}_{i\infty}$ (transformed to the loss ratio scale per $\tilde{r}_{i\infty}=\frac{\tilde{y}_{i\infty}}{p_{i}}~{}\forall i=1,...,N$). To account for uncertainty in these posterior predictions in the context of the forecasting model, we can add an error-in-variables measurement error model (Stefanski, 2000) to the forecasting models that takes the posterior means $\mathbb{E}\left[\tilde{r}_{i\infty}\right]~{}\forall i=1,...,N$ and posterior standard deviations $\mathbb{SD}\left[\tilde{r}_{i\infty}\right]~{}\forall i=1,...,N$ as input and models the “true” underlying ultimate loss ratio ${r^{\prime}}_{i\infty}$. For both the random walk and mean reversion models, this involves the following modification to the likelihood expression:
@@
In practice, there are many potential measurement error assumptions one can make, and the above is just one example. In the current context, lognormal measurement error is reasonable given that the development model predictions are generated from a lognormal distribution. An alternative approach would be to jointly model the loss development and forecasting stages such that there is no need for an intermediate summarization step and subsequent measurement error modeling (akin to Goold (2024), who jointly models the body and tail stages of loss development). However, in practice we find that it is useful to separate the loss development from forecasting stages, which allows for more intuitive setting of priors on the individual models and easier investigation of model behavior. Further, having separate stages keeps the flow of information forward such that the forecasting model does not influence development model parameter estimation.
@@ 3.3 Deriving data-driven priors
3.4 Model validation with simulated data
@@ 3.4.2 Prior and posterior predictive checks
Figure 3: Prior predictions for three random commercial auto programs from the Meyers (2015) dataset. Shaded intervals indicate 50% and 95% credible intervals, and points are true losses at $y_{i(10)}$.
@@
Figure 4: Posterior predictions for three random commercial auto programs from the Meyers (2015) dataset. The programs here are the same as in Figure 3. Shaded intervals indicate 50% and 95% credible intervals, and points are true losses at $y_{i(10)}$.
3.5 Backtesting workflow using empirical data
A model can pass all model validation checks but then fail to perform well on real-world prediction tasks given the wide range of different loss triangles that we encounter from day-to-day. Therefore, large-scale out-of-sample tests are crucial to determine if a given model (or model pipeline) is suited for production use. Further, we often have multiple competing models, where each model makes different assumptions regarding how losses change over time. For example, equation 7 assumes mean reversion toward some latent ultimate loss, whereas equation 6 assumes no such latent ultimate loss or reversion process. An added benefit of conducting out-of-sample tests is that we can take advantage of modern Bayesian stacking methods that combine predictions from competing models, allowing for a combined performance better than the best fitting individual model. In the following sections, we overview our out-of-sample backtesting workflow and our use of stacking to improve model performance.
@@ 3.5 Backtesting workflow using empirical data
For demonstration purposes, here we will rely on an open dataset described by Meyers (2015), which contains about 50 loss triangles for each of four different lines of business, including private passenger auto, worker’s compensation, commercial auto, and general liability lines. Each triangle in the Meyers dataset is 10 accident years by 10 development lags, which allows for us to fit our loss development to loss forecasting pipeline on each triangle exactly once, using the loss from the most recent accident year, development lag pair as the future, left-out datapoint for testing purposes (see Figure 5). Note that this scheme assumes $y_{i\infty}=y_{i(10)}~{}\forall i=1,...,N$, which may not be true in practice for longer-tailed lines of business. However, given data constraints for this demonstration, we believe this is a reasonable assumption. We use $y_{i\infty}$ and $y_{i(10)}$ interchangeably in sections that follow.
@@ 3.5.1 Stacking of predictive distributions
For demonstration purposes, we will use the stacking model described by Yao et al. (2018), which estimates a single weight for each model. The model can be easily fit using bayesblend – an open-source Python library that we have developed that implements various different stacking models (Haines and Goold, 2024). To fit the stacking model, we first computed the posterior log likelihood values and posterior predictions for each datapoint in the test set as illustrated in Figure 5. We did so for both the random walk and mean reversion forecasting models. Note that each posterior log likelihood in the test set is derived based on the true loss, which the models are not trained on. We then used the posterior log likelihoods and associated posterior predictions for each model as input to the MleStacking class in bayesblend, which outputs a blend of the posterior log likelihoods and predictions given the estimated model weights.
@@ 3.5.3 Model performance results
Figure 6: ELPD (left) and RMSE (right) differences (+/- 2 standard errors on the difference), ordered by Validation dataset performance for each model and line of business. The best-performing model is the reference model, shown at the top of each panel. Absolute ELPD and RMSE values are displayed above. Positive ELPD differences with an uncertainty interval that does not cross zero indicates a credible difference at the 95% level in favour of the top model. Negative RMSE differences with an uncertainty interval that does not cross zero indicates a credible difference in favor of the top model. Note that in cases where non-reference models do not have visible intervals, the uncertainty intervals are just too narrow to be visible on the plot.
@@ 3.6 From Forecasts to Cashflows
Once we have selected a model configuration that can forecast ultimate losses for future accident years to a desired level of accuracy (including calibration, RMSE, etc.), the next step is to estimate cashflows that can be used to understand the profitability of the insurance program. There are two types of cashflows that are relevant for this step: (1) the “underwriting” cashflow, which captures primarily how losses are paid out across time for the future accident year being securitized; and (2) the “investment” cashflow, which captures interest on invested premium dollars (i.e. the “float”) as well as bespoke deal structure terms (collateralization schemes, profit commission to cedants, etc.). For the current demonstration, we will focus primarily on the underwriting cashflow, as the investment cashflow is a deterministic function of underwriting cashflow (i.e. it is mostly an exercise in accounting) so is outside the scope the current work.
@@ 3.6.1 Estimating future underwriting cashflows
$\begin{split}\tilde{y}_{i(9)}^{(s)}&=\tilde{y}^{(s)}_{i(10)}/\alpha^{(s)}_{9}\\ \tilde{y}_{i(8)}^{(s)}&=\tilde{y}^{(s)}_{i(9)}/\alpha^{(s)}_{8}\quad\forall s=1,...,S\\ &\vdots\\ \tilde{y}_{i(1)}^{(s)}&=\tilde{y}^{(s)}_{i(2)}/\alpha^{(s)}_{1},\end{split}$ (16)
where $\alpha^{(s)}_{j}$ is the $s$th posterior sample of the $j$th age-to-age factor. The result is a full posterior distribution of paid loss estimates for each future accident year $i$ (see Figure 8). Across development, these paid loss estimates, relative to premiums, correspond to the underwriting cashflow associated with the future accident year.
Figure 8: Underwriting cashflows for future accident years based on historic development patters and forecasted ultimate losses for a random selection of commercial auto programs in the Meyers data. Results are based on the development and forecasting models fitted during the backtests described above. Underwriting cashflows were estimated per the walk-back method described in equation 16 using random programs selected from the Backtest workflow (see Section 3.5) illustration. Forecasted ultimate losses were based on the stacked predictions from the random walk (equation 6) and mean reversion (equation 7) models, and age-to-age factors were obtained from the chain-ladder (equation 4) and generalized Bondy (equation 5) models fitted to each program.
@@ 3.6.2 Underwriting to investment cashflows
4 Working example
In this section, we apply the full Bayesian workflow presented in Section 3 to the Swiss private liability loss development triangle from Balona and Richman (2020) (refer to Tables 23-24). These data are originally from Gisler (2009), but Balona (2020) add simulated premium data to the triangle that is needed for our workflow, so we will refer to Balona and Richman (2020) as the source from hereon out. The Swiss private liability triangle has 18 years of history spanning from 1979 to 2015, which originate from a major Swiss insurer. Note that the exact figures have been adjusted for privacy purposes–we report values directly from Balona and Richman (2020) in thousands of dollars, but emphasize that the scaling is arbitrary. As seen in Figure 9, the data exhibit the typical long-tailed, early-lag volatility that characterizes many insurance programs undergoing casualty securitization. As such, it makes for an insightful example of how to apply the Bayesian workflow in practice.
Figure 9: Observed loss ratio development patterns from the Swiss liability program in Balona and Richman (2020), indicating a generally long-tailed structure with variability in early development years.
4.1 Loss development
Following the historical loss development workflow from Section 3.1, we separate loss development into a “body” process (via the Bayesian chain-ladder model in equation 4) and a “tail” process (via the generalized Bondy model in equation 5). Because private liability falls under the more general umbrella of general liability insurance, we specify the the chain-ladder and generalized bondy training thresholds using similar values as for those used to backtest the general liability line above (see Section 3.5). Specifically, we set $\tau=6$ and $\rho_{1},\rho_{2}=(4,15)$. We also used the same prior distributions as specified in equations 4 and 5 for the chain-ladder and generalized Bondy models, respectively.
Note that the upper threshold we set for the tail model ($\rho_{2}$) is higher here because the Swiss liability triangle has 18 years of history available, whereas the Meyers data used for backtests only had 10 years of history. Further, unlike for the backtest, we do not set the upper tail threshold to the maximum possible given the training data (i.e. $\rho_{2}=18$). Including the full training data means that tail development factors for the most recent development lags are informed only by the most historic accident years where data are available. By not fitting the tail model to the most recent development lags, we allow for the model to be informed only by development patterns where we have multiple accident years, which can make the tail estimates more robust to noise that may otherwise be present in 1 or 2 early accident years.
Figure 10 shows posterior predictions from the development models on the Swiss private liability program, faceted by accident year to better visualize the uncertainty in the predictions. Generally, we see that the model predictions are mostly flat after 6-7 years of development. It is also clear that the model is more uncertain for more recent accident years, which is expected given we only have one or more observed years of development for the latest accident years. Overall, the importance of the development model is most readily observed for later accident years, where the model captures the expected increase in cumulative losses across early development periods before converging toward the ultimate loss ratios for each year.
Figure 10: Observed loss ratio development patterns along with posterior predictions from loss development models for the Swiss liability program in Balona and Richman (2020). The predicted ultimate loss ratio means and standard deviations (i.e. predicted means and standard deviations at development year 30) are used as input to the forecasting model described in Section 4.2. The shaded regions correspond to the 50% and 95% credible intervals for the predictions.
4.2 Ultimate loss forecasting
Given the predicted ultimate losses obtained for historic accident years (see Figure 10), the next step is to predict ultimates for future accident years. For the current example, we will use the same forecasting models as described in Section 3.2. Specifically, we fit both the random walk (equation 6) and mean reversion (equation 7) models, both with the measurement error specification in equation 8. After generating one-year-ahead predictions from each fitted model, we then blend together the posterior predictive distributions to arrive at the final one-year-ahead ultimate loss ratio prediction.
For blending the forecasting model predictions, we used the same MleStacking model as in our backtests (see Section 3.5.1). Recall that in our backtests, we trained a separate stacking model for each of the four lines of business in the Meyers (2015) dataset. To blend the Swiss liability triangle predictions, we used the stacking model trained on the general liability subset of the Meyers data. The MleStacking model only assigns a single static weight to each forecasting model, which in our case produced weights of 0.55 for the random walk model and 0.45 for the mean reversion model. Given these weights, the blended ultimate loss ratio distribution is generated by sampling the posterior predictive distributions from each forecasting model in proportion to each model’s respective weight.
Finally, for our forecasting model priors, we used data-driven priors as described in Section 3.3. Specifically, when conducting the backtests for each line of business in the Meyers dataset (see Section 3.5), we saved the posterior distributions from the hierarchically estimated forecasting models (see equation 10), summarized the group-level parameters for each line of business, and used these summaries as the priors for each forecasting model fitted to the Swiss liability triangle:
$\eta_{\text{init}}$ $\sim\mathcal{N}(-0.92,0.41)$ (Random Walk)
$\epsilon$ $\sim\mathcal{N}(-2.17,0.22)$
$\gamma_{1}$ $\sim\mathcal{N}(-2.18,0.96)$
$\gamma_{2}$ $\sim\mathcal{N}(-2.13,0.92)$
$\eta_{\text{init}}$ $\sim\mathcal{N}(-1.49,0.86)$ (Mean Reversion)
$\epsilon$ $\sim\mathcal{N}(-2.66,0.15)$
$\mu$ $\sim\mathcal{N}(-0.65,0.52)$
$\phi$ $\sim\mathcal{N}(1.27,0.95)$
$\gamma_{1}$ $\sim\mathcal{N}(-2.23,0.91)$
$\gamma_{2}$ $\sim\mathcal{N}(-2.24,0.86)$
Figure 11 shows the posterior predictive distributions for each forecasting model across accident years, including the one-year-ahead predictions. The models capture the same general patterns, although the mean reversion model tends to produce predictions with less uncertainty than the random walk model. This occurs because the mean reversion process pulls year-to-year predictions toward the mean of the historical data, which reduces overall variability in the predictions. Further, the mean reversion model produces a slightly higher ultimate loss ratio mean than the random walk model. The difference in both the expectation and uncertainty of predictions across models highlights the importance of model blending, allowing us to combine models as opposed to selecting just one.
Figure 11: Posterior predictive distributions for each forecasting model across accident years for the Swiss liability triangle from Balona and Richman (2020). The shaded regions correspond to the 50% and 95% credible intervals for the predictions.
4.3 Underwriting Cashflows
The final step we will cover in our working example is to predict one-year-ahead underwriting cashflows given the forecasted ultimate loss ratio predictions and the historic loss development patterns. Here, we use the walk-back method outlined in equation 16. As explained in Section 3.6.1, the walk-back method starts from the one-year-ahead ultimate loss ratio obtained from the forecasting models, then uses the age-to-age factors from the loss development model (Section 4.1 above) to iteratively predict back to the first development lag. This process results in a full time-series of posterior predicted cashflows for the one-year-ahead accident year, which represent the paid losses expected to emerge for the future accident year across time.
Figure 12 shows the predicted underwriting cashflows for the Swiss liability triangle. It is clear from Figure 12 that paid losses are expected to converge to the ultimate loss after approximately 6 years, with an expected loss ratio of 56% and a 97.5%ile loss ratio of about 72%. Given the relatively low expected ultimate loss ratio along with the development pattern, this program is a good candidate for securitization (see Section 3.6.2).
Figure 12: Predicted paid loss development for the one-year-ahead accident year for the Swiss liability triangle from Balona and Richman (2020). Paid losses for future accident years represent the expected underwriting cashflows associated with the ILS deal. The shaded regions correspond to the 50% and 95% credible intervals for the predictions.